# MSDNet: Multi-Scale Decoder for Few-Shot Semantic Segmentation via Transformer-Guided Prototyping

### v1

@@ Abstract
Few-shot Semantic Segmentation addresses the challenge of segmenting objects in query images with only a handful of annotated examples. However, many previous state-of-the-art methods either have to discard intricate local semantic features or suffer from high computational complexity. To address these challenges, we propose a new Few-shot Semantic Segmentation framework based on the transformer architecture. Our approach introduces the spatial transformer decoder and the contextual mask generation module to improve the relational understanding between support and query images. Moreover, we introduce a multi-scale decoder to refine the segmentation mask by incorporating features from different resolutions in a hierarchical manner. Additionally, our approach integrates global features from intermediate encoder stages to improve contextual understanding, while maintaining a lightweight structure to reduce complexity. This balance between performance and efficiency enables our method to achieve state-of-the-art results on benchmark datasets such as $PASCAL-5^{i}$ and $COCO-20^{i}$ in both 1-shot and 5-shot settings. Notably, our model with only 1.5 million parameters demonstrates competitive performance while overcoming limitations of existing methodologies. https://github.com/amirrezafateh/MSDNet
Index Terms:
@@ I Introduction
Semantic segmentation is a key task in computer vision, where each pixel of an image is labeled as part of a specific category. This is important in many areas like OCR, autonomous driving, medical imaging, and scene understanding [1, 2, 3]. To perform this task well, models need to learn detailed object boundaries. In recent years, deep Convolutional Neural Networks (CNNs) have made big improvements in this area [4]. However, these high-performing models usually need large datasets with lots of labeled examples [5, 6], which takes a lot of time and effort to create. In real-world scenarios, like in medical imaging or other fields where labeled data is limited, this becomes a big problem [7, 8]. To solve this, Few-shot Semantic Segmentation (FSS) has become a useful approach.
@@
FSS tries to segment new object classes in images using only a few labeled examples, called support images, that show the target class [9]. This method helps reduce the need for large datasets, making it more practical for real-world use [10]. Addressing the challenges of FSS requires handling differences in texture or appearance between the target object in the query image and similar objects depicted in the support examples. Effectively using the relationship between the query image and the support examples is essential in tackling FSS.
@@
FSS can be widely categorized into two groups: Prototype-based approaches and Pixel-wise methods. As shown in Figure 1-LABEL:sub@fig:1_a, prototype-based approaches involve abstracting semantic features of the target class from support images through a shared backbone network [11]. This process results in feature vectors called class-wise prototypes, which are obtained using techniques such as class-wise average pooling or clustering. These prototypes are then combined with query features through operations like element-wise summation or channel-wise concatenation. The combined features are refined by a decoder module to classify each pixel as either the target class or background [12]. In contrast, as shown in Figure 1-LABEL:sub@fig:1_b, pixel-wise methods take a different approach by focusing directly on pixel-level information rather than compressing it into prototypes. These methods aim to predict the target class for each pixel in the query image by comparing it directly with corresponding pixels in the support images. To achieve this, they establish pixel-to-pixel correlations between the support and query features, which allows the model to find precise matches even when the object’s appearance varies [13]. This process is often enhanced by attention mechanisms, like those found in Transformer models, which help the model focus on important relationships between pixels. By avoiding the need for prototypes, Pixel-wise methods aim to preserve more detailed information, allowing for finer-grained segmentation [14, 15].
While both groups have demonstrated efficacy, they also have certain limitations. Prototype-based methods may inadvertently discard complex local semantic features specific to the target class in support images. This can lead to coarse segmentation of the target class in query images, especially for objects with complex appearances. On the other hand, while pixel-wise methods have notably improved performance compared to prototype-based approaches, they grapple with computational complexity due to dot-product attention calculations across all pixels of support features and query features. Moreover, a large amount of pixel-wise support information can lead to confusion in attention mechanisms [13]. Also, a shared limitation across both approaches is the lack of use of encoder middle features in the decoder section. Many methods in both categories employ straightforward decoders that fail to incorporate encoder middle features. However, in few-shot scenarios where data samples are limited, leveraging the global features captured by the encoder in the decoder phase can prove to be highly beneficial.
@@
Inspired by recent developments, we aim to develop a straightforward and effective framework to address limitations in FSS methods. A notable approach gaining traction is the Query-based [^1] For differentiating it from the conventional term ”query” frequently employed in FSS, we capitalize ”Query” when referring to the query sequence within the Transformer architecture. Transformer architecture, which has demonstrated versatility across various computer vision tasks, including few-shot learning scenarios [16, 17]. This architecture utilizes learnable Query embeddings derived from support prototypes, enabling nuanced analysis of their relationships within the query feature map.
@@
Inspired by previous works, as shown in Figure 1-LABEL:sub@fig:1_c, we have designed a novel Transformer-based module, known as the Spatial Transformer Decoder (STD), to enhance the relational understanding between support images and the query image. This module operates concurrently with the multi-scale decoder. Within the STD module, we introduce a common strategy: Using the prototype of support images as a Query, while utilizing the features extracted from the query image as both Value and Key embeddings inputted into the Transformer decoder. This formulation allows the Query to effectively focus on the semantic features of the target class within the query image. Furthermore, to reduce the impact of information loss resulting from the abstraction of support images into a feature vector named the ’support prototype,’ we integrate global features from the intermediate stages of the encoder, which are fed with the support images, into our decoder. Incorporating these features allows us to leverage features from different stages of the encoder, thereby enriching the decoder’s contextual understanding. Additionally, we introduce the Contextual Mask Generation Module (CMGM) to further augment the model’s relational understanding, operating alongside the STD and enhancing the model’s capacity to capture relevant contextual information.
II Related Works
II-A Semantic Segmentation
@@
Semantic segmentation, a crucial task in computer vision, involves labeling each pixel in an image with a corresponding class [18, 19]. CNNs significantly advanced semantic segmentation by replacing fully connected layers with convolutional layers, enabling the processing of images of various sizes [20]. Since then, subsequent advancements have focused on enhancing the receptive field and aggregating long-range context in feature maps. Techniques such as dilated convolutions [21], spatial pyramid pooling [22], and non-local blocks [23] have been employed to capture contextual information at multiple scales. More recently, Transformer-based backbones, including SegFormer [24], Segmenter [25], and SETR [26], have been introduced to better capture long-range context in semantic segmentation tasks. Further enhancing this approach, hierarchical architectures like the Swin Transformer [27] have achieved state-of-the-art performance by using shifted windows in their general-purpose backbones. In parallel, self-supervised pretraining strategies, such as the masked image modeling used in BEiT [28], have also shown strong results, fine-tuning directly on the semantic segmentation task and pushing the boundaries of model performance.
@@
Semantic segmentation tasks typically involve per-pixel classification. as demonstrated by approaches like MaskFormer [29] and Mask2Former [30], which predict binary masks corresponding to individual class labels. Older architectures, such as UNet [31], PSPNet [32], and Deeplab [33, 34], have also significantly contributed to the field by incorporating features like global and local context aggregation and dilated convolutions to increase the receptive field without reducing resolution. Building upon these foundational approaches, more recent studies, including CRGNet [35] and SAM [36], have focused on further improving model performance, exploring new techniques to enhance accuracy in segmentation tasks. Despite the progress made in per-pixel classification, addressing the challenge of segmenting unseen classes remains an open area for future research
II-B Few-Shot Semantic Segmentation
@@
FSS is a challenging task in computer vision, wherein the objective is to segment images with limited annotated examples, known as support images. Approaches to FSS can be categorized into various groups based on their primary aims and methodologies employed [37]. One significant challenge in FSS is addressing the imbalance in details between support and query images. Methods like PGNet [38] and PANet [39] aim to eliminate inconsistent regions between support and query images by associating each query pixel with relevant parts of the support image or by regularizing the network to ensure its success regardless of the roles of support and query. But methods like ASGNet [37], on the other hand, focuses on finding an adaptive quantity of prototypes and their spatial expanses determined by image content, utilizing a boundary-conscious superpixel algorithm.
@@
Another critical aspect of FSS is bridging the inter-class gap between base and novel datasets. Approaches like RePRI [40] and CWT [41] address this gap by fine-tuning over support images or episodically training self-attention blocks to adapt classifier weights during both training and testing phases. Additionally, architectures designed for supervised learning often trouble recognizing objects at different scales in few-shot scenarios. To address this issue, new methods have been developed to allow information exchange between different resolutions [42].
@@
Moreover, ensuring the reliability of correlations between support and query images is essential in FSS. Methods like HSNet [43] and CyCTR [44] utilize attention mechanisms to filter out erroneous support features and focus on beneficial information. VAT [45], meanwhile, employs a cost aggregation network to aggregate information between query and support features, leveraging a high-dimensional Swin Transformer to impart local context to all pixels.
@@
III Proposed method
III-A Problem Definition
@@
III-B Overview
Figure 2: the overview of the proposed method
Given a support set $S=\left\{I_{s}^{i},M_{s}^{i}\right\}$ and a query image $I_{q}$, the objective is to generate the binary segmentation mask for $I_{q}$, identifying the same class as the support examples. To address this task, we introduce a straightforward yet robust framework, outlined in Figure 2. For simplicity, we illustrate a 1-shot setting within the framework, but this can be easily generalized to a 5-shot setting as well. The proposed method comprises several key components, including a shared pretrained backbone, support prototype, CMGM, a multi-scale decoder, and STD. These elements collectively contribute to the model’s ability to accurately segment objects of interest in the query image based on contextual information provided by the support set. In the following, we’ll take a closer look at each component, explaining its role and how it interacts within our framework.
III-B1 Backbone
In our proposed framework, we adopt a modified ResNet architecture, initially pre-trained on the ImageNet dataset, to serve as the backbone for feature extraction from raw input images, ensuring that the size of the output of each block does not reduce below a specified dimension. For instance, like [46], we define that the output sizes from $conv2\_x$ to $conv5\_x$ are maintained at $60\times 60$ pixels. Specifically, we utilize a ResNet with shared weights between support and query images. This type of ResNet maintains the spatial resolution of feature maps at $60\times 60$ pixels from the $conv2\_x$ stage forward, preserving finer details crucial for accurate segmentation. We extract high-level features ($conv5\_x$), as well as mid-level features ($conv3\_x$ and $conv4\_x$) from both support and query images using the backbone.
The mid-level features of the support image are denoted as $X_{s}^{conv3}$ and $X_{s}^{conv4}$, while the high-level features are denoted as $X_{s}^{conv5}$. Similarly, for the query image, the mid-level features are represented as $X_{q}^{conv3}$ and $X_{q}^{conv4}$, and the high-level features as $X_{q}^{conv5}$. To integrate mid-level features across different stages, we concatenate the mid-level feature maps from $conv3\_x$ and $conv4\_x$ stages and apply a $1\times 1$ convolution layer to yield a merged mid-level feature map, denoted as $X_{s}^{merged}$. This merging process ensures that the resultant feature map retains essential information from both mid-level stages, enhancing the model’s ability to capture diverse contextual information (Equation 1, Equation 2).
@@ III-B2 Support Prototype
III-B3 Contextual Mask Generation Module (CMGM)
The CMGM is a novel component introduced by our framework, designed to enhance the contextual understanding between support and query images in FSS tasks. As shown in Figure 3, CMGM leverages the feature representations extracted from both the support and query images to generate a contextual mask that encapsulates pixel-wise relations indicative of the target object. This process involves computing the cosine similarity between the query feature vector and the support feature vector. Mathematically, cosine similarity $cos(q,s)$ is calculated as the dot product of the normalized query and support feature vectors. In a five-shot scenario, where there are five support examples, five cosine similarities are computed and subsequently averaged, yielding a novel cosine similarity measure representative of the collective support set.
Figure 3: the overview of CMGM
III-B4 Multi Scale Decoder

### v3

@@ Abstract
: Few-shot Semantic Segmentation addresses the challenge of segmenting objects in query images with only a handful of annotated examples. However, many previous state-of-the-art methods either have to discard intricate local semantic features or suffer from high computational complexity. To address these challenges, we propose a new Few-shot Semantic Segmentation framework based on the Transformer architecture. Our approach introduces the spatial transformer decoder and the contextual mask generation module to improve the relational understanding between support and query images. Moreover, we introduce a multi scale decoder to refine the segmentation mask by incorporating features from different resolutions in a hierarchical manner. Additionally, our approach integrates global features from intermediate encoder stages to improve contextual understanding, while maintaining a lightweight structure to reduce complexity. This balance between performance and efficiency enables our method to achieve competitive results on benchmark datasets such as $PASCAL\text{-}5^{i}$ and $COCO\text{-}20^{i}$ in both 1-shot and 5-shot settings. Notably, our model with only 1.5 million parameters demonstrates competitive performance while overcoming limitations of existing methodologies. https://github.com/amirrezafateh/MSDNet
Keywords:
@@ 1 Introduction
Semantic segmentation is a key task in computer vision, where each pixel of an image is labeled as part of a specific category. This is important in many areas like autonomous driving, medical imaging, and scene understanding zhang2024segment . To perform this task well, models need to learn detailed object boundaries. In recent years, deep Convolutional Neural Networks (CNNs) have made big improvements in this area sun2024remax . However, these high-performing models usually need large datasets with lots of labeled examples bi2024prompt ; bi2024agmtr ; bi2023not , which takes a lot of time and effort to create. In real-world scenarios, like in medical imaging or other fields where labeled data is limited, this becomes a big problem. To solve this, Few-shot Semantic Segmentation (FSS) has become a useful approach.
@@
FSS tries to segment new object classes in images using only a few labeled examples, called support images, that show the target class askari2025enhancing ; mao2022bidirectional ; mao2024body . This method helps reduce the need for large datasets, making it more practical for real-world use. Addressing the challenges of FSS requires handling differences in texture or appearance between the target object in the query image and similar objects depicted in the support examples. Effectively using the relationship between the query image and the support examples is essential in tackling FSS. FSS can be widely categorized into two groups: Prototype-based approaches and Pixel-wise methods.
@@
Prototype-based approaches involve abstracting semantic features of the target class from support images through a shared backbone network. This process results in feature vectors called class-wise prototypes, which are obtained using techniques such as class-wise average pooling or clustering. These prototypes are then combined with query features through operations like element-wise summation or channel-wise concatenation. The combined features are refined by a decoder module to classify each pixel as either the target class or background li2024dual ; sun2023attentional ; liu2023learning ; ding2023self (Figure 1-1(a)). Pixel-wise methods take a different approach by focusing directly on pixel-level information rather than compressing it into prototypes. These methods aim to predict the target class for each pixel in the query image by comparing it directly with corresponding pixels in the support images. To achieve this, they establish pixel-to-pixel correlations between the support and query features, which allows the model to find precise matches even when the object’s appearance varies. This process is often enhanced by attention mechanisms, like those found in Transformer models, which help the model focus on important relationships between pixels. By avoiding the need for prototypes, Pixel-wise methods aim to preserve more detailed information, allowing for finer-grained segmentation xu2023self ; kang2023distilling ; shi2022dense . An example of this is illustrated in Figure 1-1(b).
While both prototype-based and pixel-wise approaches have demonstrated efficacy in few-shot semantic segmentation, they also exhibit key limitations. Prototype-based methods often compress the semantic features of the support images into a single vector, potentially discarding fine-grained spatial information necessary for accurate segmentation—especially for complex object classes. Pixel-wise methods address this by directly comparing individual pixels across support and query images, but they suffer from high computational costs due to full dot-product attention and can become unstable when overloaded with dense pixel-wise support features xu2023self . A common limitation shared by both approaches is the under utilization of intermediate encoder features during decoding. Most methods rely on shallow or single-scale decoders that do not effectively incorporate mid-level representations from the encoder, missing valuable contextual information. This is particularly problematic in few-shot settings, where richer features are essential for generalizing from limited samples. These challenges highlight a clear gap: the need for a lightweight yet semantically expressive framework that effectively captures both relational understanding and multi-scale context for robust few-shot segmentation.
@@
Inspired by recent developments, we aim to develop a straightforward and effective framework to address limitations in FSS methods. A notable approach gaining traction is the Query-based [^1] For differentiating it from the conventional term ”query” frequently employed in FSS, we capitalize ”Query” when referring to the query sequence within the Transformer architecture. Transformer architecture, which has demonstrated versatility across various computer vision tasks, including few-shot learning scenarios su2024roformer ; tian2024survey ; luo2023closer ; cao2022prototype . This architecture utilizes learnable Query embeddings derived from support prototypes, enabling nuanced analysis of their relationships within the query feature map.
@@
Inspired by previous works, we have designed a novel Transformer-based module, known as the Spatial Transformer Decoder (STD), to enhance the relational understanding between support images and the query image. This module operates concurrently with the multi-scale decoder. The core architecture of our approach is shown in Figure 1-1(c). Within the STD module, we introduce a common strategy: Using the prototype of support images as a Query, while utilizing the features extracted from the query image as both Value and Key embeddings inputted into the Transformer decoder. This formulation allows the Query to effectively focus on the semantic features of the target class within the query image. Furthermore, to reduce the impact of information loss resulting from the abstraction of support images into a feature vector named the ’support prototype,’ we integrate global features from the intermediate stages of the encoder, which are fed with the support images, into our decoder. Incorporating these features allows us to leverage features from different stages of the encoder, thereby enriching the decoder’s contextual understanding. Additionally, we introduce the Contextual Mask Generation Module (CMGM) to further augment the model’s relational understanding (not shown in Figure 1-1(c)), operating alongside the STD and enhancing the model’s capacity to capture relevant contextual information.
In summary, our contributions include:
  1. 1.
    We propose MSDNet, a novel and lightweight framework for few-shot semantic segmentation, which incorporates a STD. In contrast to conventional designs, our STD uses the support prototype as the Query and the query feature map as the Key and Value in a multi-head cross-attention mechanism, enhancing semantic alignment between support and query features. Despite having only 1.5M learnable parameters, our model achieves competitive performance on standard benchmarks.
  2. 2.
    We introduce a multi-scale decoder architecture that hierarchically refines segmentation masks using progressively integrated mid-level and high-level support features. This approach differs from most prior FSS methods, which commonly rely on shallow or single-scale decoders, and enables more precise mask generation with spatial detail.
  3. 3.
    We develop a novel CMGM, which enhances pixel-wise relational understanding by computing cosine similarities between support and query features. This module provides a semantic prior that guides subsequent processing stages more effectively than traditional feature concatenation or simple prototype averaging.
  4. 4.
    We conduct comprehensive evaluations on the $PASCAL\text{-}5^{i}$ and $COCO\text{-}20^{i}$ benchmarks in both 1-shot and 5-shot settings. Our model consistently ranks among the top-performing methods across all folds, confirming its effectiveness and efficiency in a variety of segmentation scenarios.
2 Related Works
2.1 Semantic Segmentation
@@
Semantic segmentation, a crucial task in computer vision, involves labeling each pixel in an image with a corresponding class rizzoli2024source ; rezvani2024single ; zhou2024cross . CNNs significantly advanced semantic segmentation by replacing fully connected layers with convolutional layers, enabling the processing of images of various sizes saber2024efficient ; rezvani2025fusionlungnet . Since then, subsequent advancements have focused on enhancing the receptive field and aggregating long-range context in feature maps. Techniques such as dilated convolutions yu2015multi , spatial pyramid pooling he2015spatial , and non-local blocks wang2018non have been employed to capture contextual information at multiple scales. More recently, Transformer-based backbones, including SegFormer xie2021segformer , Segmenter strudel2021segmenter , and SETR zheng2021rethinking , have been introduced to better capture long-range context in semantic segmentation tasks. Hierarchical architectures like Swin Transformer liu2021swin have achieved SOTA performance by computing shifted windows for general-purpose backbones. Additionally, self-supervised pretraining approaches, such as masked image modeling in BEiT bao2021beit , have demonstrated competitive results by fine-tuning directly on the semantic segmentation task.
@@
Semantic segmentation tasks typically involve per-pixel classification. as demonstrated by approaches like MaskFormer cheng2021per and Mask2Former cheng2022masked , which predict binary masks corresponding to individual class labels. Older architectures, such as UNet ronneberger2015u , PSPNet zhao2017pyramid , and Deeplab chen2017deeplab ; chen2018encoder , have also significantly contributed to the field by incorporating features like global and local context aggregation and dilated convolutions to increase the receptive field without reducing resolution. Recent studies have sought to improve model performance and contribute to the advancement of semantic segmentation, with notable approaches including CRGNet xu2022consistency , and SAM kirillov2023segment , among others. While significant progress has been made in understanding and classifying images at the pixel level, further advancements are needed to effectively address the challenge of unseen classes in semantic segmentation.
Figure 2: the overview of the proposed method
2.2 Few-Shot Semantic Segmentation
@@
FSS is a challenging task in computer vision, wherein the objective is to segment images with limited annotated examples, known as support images. Approaches to FSS can be categorized into various groups based on their primary aims and methodologies employed li2021adaptive ; lu2021simpler ; xie2021scale ; hong2022cost . One significant challenge in FSS is addressing the imbalance in details between support and query images. Methods like PGNet zhang2019pyramid and PANet wang2019panet aim to eliminate inconsistent regions between support and query images by associating each query pixel with relevant parts of the support image or by regularizing the network to ensure its success regardless of the roles of support and query. ASGNet li2021adaptive , on the other hand, focuses on finding an adaptive quantity of prototypes and their spatial expanses determined by image content, utilizing a boundary-conscious superpixel algorithm.
@@
Another critical aspect of FSS is bridging the inter-class gap between base and novel datasets. Approaches like RePRI boudiaf2021few and CWT lu2021simpler address this gap by fine-tuning over support images or episodically training self-attention blocks to adapt classifier weights during both training and testing phases. Additionally, architectures designed for supervised learning often trouble recognizing objects at different scales in few-shot scenarios. To address this issue, new methods have been developed to allow information exchange between different resolutions kayabacsi2023elimination ; xin2022multilevel .
@@
Moreover, ensuring the reliability of correlations between support and query images is essential in FSS. Methods like HSNet min2021hypercorrelation and CyCTR zhang2021few utilize attention mechanisms to filter out erroneous support features and focus on beneficial information. VAT hong2022cost , meanwhile, employs a cost aggregation network to aggregate information between query and support features, leveraging a high-dimensional Swin Transformer to impart local context to all pixels.
@@
3 Proposed method
3.1 Problem Definition
@@
3.2 Overview
Given a support set $S=\left\{I_{s}^{i},M_{s}^{i}\right\}$ and a query image $I_{q}$, the objective is to generate the binary segmentation mask for $I_{q}$, identifying the same class as the support examples. To address this task, we introduce a straightforward yet robust framework, outlined in Figure 2. For simplicity, we illustrate a 1-shot setting within the framework, but this can be easily generalized to a 5-shot setting as well. The proposed method comprises several key components, including a shared pretrained backbone, support prototype, Contextual Mask Generation Module (CMGM), a multi-scale decoder, and Spatial Transformer Decoder (STD). These elements collectively contribute to the model’s ability to accurately segment objects of interest in the query image based on contextual information provided by the support set. In the following, we’ll take a closer look at each component, explaining its role and how it interacts within our framework.
3.2.1 Backbone
In our proposed framework, we adopt a modified ResNet architecture, initially pre-trained on the ImageNet dataset, to serve as the backbone for feature extraction from raw input images, ensuring that the size of the output of each block does not reduce below a specified dimension. For instance, like cao2022prototype , we define that the output sizes from conv2_x to conv5_x are maintained at 60x60 pixels. Specifically, we utilize a ResNet with shared weights between support and query images. This type of ResNet maintains the spatial resolution of feature maps at $60\times 60$ pixels from the conv2_x stage forward, preserving finer details crucial for accurate segmentation. We extract high-level features (conv5_x), as well as mid-level features (conv3_x and conv4_x) from both support and query images using the backbone.
The mid-level features of the support image are denoted as $X_{s}^{conv3}$ and $X_{s}^{conv4}$, while the high-level features are denoted as $X_{s}^{conv5}$. Similarly, for the query image, the mid-level features are represented as $X_{q}^{conv3}$ and $X_{q}^{conv4}$, and the high-level features as $X_{q}^{conv5}$. To integrate mid-level features across different stages, we concatenate the mid-level feature maps from conv3_x and conv4_x stages and apply a $1\times 1$ convolution layer to yield a merged mid-level feature map, denoted as $X_{s}^{merged}$. This merging process ensures that the resultant feature map retains essential information from both mid-level stages, enhancing the model’s ability to capture diverse contextual information (Equation 1, Equation 2).
@@ 3.2.2 Support Prototype
Figure 3: Spatial Transformer Decoder
3.2.3 Contextual Mask Generation Module (CMGM)
The CMGM is a novel component introduced by our framework, designed to enhance the contextual understanding between support and query images in FSS tasks. At its core, CMGM leverages the feature representations extracted from both the support and query images to generate a contextual mask that encapsulates pixel-wise relations indicative of the target object. This process involves computing the cosine similarity between the query feature vector and the support feature vector. Mathematically, cosine similarity $cos(q,s)$ is calculated as the dot product of the normalized query and support feature vectors. In a five-shot scenario, where there are five support examples, five cosine similarities are computed and subsequently averaged, yielding a novel cosine similarity measure representative of the collective support set.
The contextual mask produced by CMGM plays a foundational role in guiding the downstream decoder modules. By emphasizing pixel-wise correspondences between the support and query images, CMGM effectively filters the relevant foreground regions. This contextual guidance becomes especially important for the subsequent modules, as it narrows their focus to semantically important regions, allowing them to operate more efficiently and precisely.
3.2.4 Multi Scale Decoder